OpenStudy (anonymous):

At Water World the admission fee is $17.00, and you can rent an inner tube for $0.75 per hour. You can use the formula P = 17.00 + 0.75h to find the total cost for admission plus renting an inner tube. a. Solve the formula P = 17.00 + 0.75h for h. Explain your steps for solving the formula for h. b. For how many hours can you rent an inner tube if you only have $20.75 to spend? Explain your method for finding the answer.

OpenStudy (hoblos):

a. P = 17.00 + 0.75h 0.75h = p-17 h= (p-17)/0.75 b.p=20.75 h= (20.75-17.00)/0.75 = 5hrs

OpenStudy (anonymous):

so that is the whole answer 4 A.? this that my method?

OpenStudy (hoblos):

yup thats it

OpenStudy (anonymous):

okay thx...what is b

OpenStudy (hoblos):

ohhhh i miss understood you!! this is the answer for both parts a&b

OpenStudy (anonymous):

oh okay thanks..happy valentines day

OpenStudy (hoblos):

its my pleasure XD happy valentines day ;)